OpenStudy (anonymous):

Alright, so I need to solve these with elimination or substitution, and I am having a hard time. 2x+y=25 3y=2x-13 and y-2x=-8 4y+13x=10 Thank you in advance if you help. ;u;

OpenStudy (shubhamsrg):

you're welcome in advance as am going to help! ;)

OpenStudy (shubhamsrg):

well calculate y from eqn 1 and eqn 2.. equate them.. you'll find x thus.. same with second ques.. if you didnt get it please tell me..

OpenStudy (anonymous):

I'm not sure I'm understanding. I work better when I can see the steps of a problem.

OpenStudy (shubhamsrg):

in first ques.. from eqn 1,, y= 25 -2x from eqn 2,, y=(2x-13)/3 thus 25-2x = (2x-13)/3 calculate x from here..put that in any of the eqns and solve for y this is substitution method for elimination,,multiply both sides of 1st eqn by 3 and subtract it from 2nd .. you'll get this 6x + 3y -3y = 75 -2x +13 => 6x = 75 - 2x +13 x= ? now if this is clear can you show me 2nd ques yourself

OpenStudy (anonymous):

For the first question, I am using Substitution method: 2x + y = 25 -------- eqn (1) 3y = 2x - 13--------eqn (2) First re-arrange eqn (2) to look like eqn (1) 2x - 3y = 13 ------eqn (3) Make y the subject for eqn (1) y = 25 - 2x --------eqn (4) Put the answer of y in eqn (4) into anywhere you see y in eqn (3) 2x - 3(25 - 2x) = 13 Expand the bracket and solve for x 2x - 75 + 6x = 13 Group like-terms on one side and solve for x 2x + 6x = 13 + 75 8x = 88 Divide both sides by 8 to make x stand alone. \[x = \frac{88}{8}\] x = 11 Put the answer of x (x = 11) into any equation above and solve for y. I am choosing eqn(3) 2x - 3y = 13 ------eqn (3) 2(11) - 3y = 13 22 - 3y = 13 -3y = 13 - 22 -3y = -9 Divide both sides by -3 to allow y to standalone \[y = \frac{-9}{-3}\] y = 3 Hence, x = 11 and y = 3 Follow the same procedure and solve the second question.

OpenStudy (anonymous):

Thanks alot! This helps! :D

OpenStudy (anonymous):

You welcome